### v1

# Family Unification in a Six Dimensional Theory with an Orthogonal Gauge Group

NITEP 242
Family Unification in a Six Dimensional Theory
with an Orthogonal Gauge Group
Nobuhito Marua,b and Ryujiro Nagoa
aDepartment of Physics, Osaka Metropolitan University,
Osaka 558-8585, Japan
bNambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP),
Osaka Metropolitan University, Osaka 558-8585, Japan
We propose a simple model of family unification, which is a six dimensional $SO(20)$ gauge theory with a single fermion in the spinorial representation. After compactification to five dimensions, our model gives a five dimensional model where the Standard Model Higgs field is unified into the fifth component of the five dimensional gauge field as well as three generations of quarks and leptons are unified into a single spinor field.
The origin of three generations of quarks and leptons is one of the mysteries in the Standard Model (SM). An idea to solve this problem is to extend the SM gauge group or the grand unified gauge group to a larger symmetry group $G$ incorporating the structure of three generations. The three generations of SM fermions are embedded into ideally a single fermion or non-repetitive set of fermions in some irreducible representations of $G$. After symmetry breaking of the group $G$ to the SM group or the grand unified group, three generation of the SM fermions are obtained as massless fields. This approach is often called as “Family unification”. There have been many proposals for the family unification from the various viewpoints, so far [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
In our previous paper [25], we have proposed a six dimensional (6D) $SU(14)$ model of family unification based on the grand gauge-Higgs unification (GGHU) where three generation quarks and leptons are obtained from several fermions in totally anti-symmetric representations. A remarkable point in this model is that the SM gauge fields and Higgs field are unified in the higher dimensional gauge field as well as the the unification of three generations of the SM fermions. However, there also remained some unsatisfactory points. First, several fermions were necessary for obtaining three generations. As mentioned above, it is desirable to obtain three generations from a single fermion. Second, some unwanted massless fermions not the SM fermions were often left after symmetry breaking. To provide masses for them, we have to introduce some additional interactions and fields, then the model becomes complicated.
In this letter, we propose a new simple model of family unification avoiding the above unsatisfactory points. This model is also based on the GGHU, but an orthogonal group is employed as a gauge group. As was discussed in [25], the reason of considering the orthogonal group is that a spinor representation in the orthogonal group has a very large dimension with respect to the size of the gauge group. In fact, three generation of the SM fermions can be obtained from a single fermion belonging to the spinorial representation by considering some simple dynamical assumptions. Our strategy of the model building is as follows. We start from a 6D SO gauge theory with a single fermion in a spinorial representation. By compactifying the sixth spatial dimension on an orbifold $S^{1}/Z_{2}$, we obtain a model of 5D gauge-Higgs unification incorporating three generation structure.
Let us consider a 6D $SO(20)$ gauge theory with a single fermion in a spinorial representation ${\bf 1024}$, which is compactified on an orbifold $S^{1}/Z_{2}$ in a sixth dimension with a radius $R_{6}$.
First, the $Z_{2}$ parity transformations for the gauge field are given as follows.
$A_{M}(x_{\mu},x_{5},x_{i}-x_{6})$ $=P_{i}A_{M}(x_{\mu},x_{5},x_{i}+x_{6})P_{i}^{\dagger}~{}(M=0,1,2,3,5),$ (1)
$A_{6}(x_{\mu},x_{5},x_{i}-x_{6})$ $=-P_{i}A_{6}(x_{\mu},x_{5},x_{i}+x_{6})P_{i}^{\dagger},$ (2)
where $P_{i}~{}(i=0,1)$ are parity matrices around the fixed points $x_{0}=0,x_{1}=\pi R_{6}$, which are chosen so that the desirable gauge symmetry may be remained.
$SO(20)$ generators are represented as follows.
$\sigma_{2}\otimes S_{10},\quad\sigma_{0}\otimes A_{10},\quad\sigma_{2}\otimes I_{10},\quad\sigma_{1}\otimes A_{10},\quad\sigma_{3}\otimes A_{10},$ (3)
where $\sigma_{i}(i=1,2,3)$ are Pauli matrices, $S_{N}(A_{N})$ is $N\times N$ (anti-)symmetric matrices, respectively. If we choose a $Z_{2}$ parity matrix at $x_{6}=0$ as
$P_{0}=\sigma_{0}\otimes I_{6,4},$ (4)
where $\sigma_{0}=\rm{diag}(1,1)$ and $I_{6,4}\equiv\rm{diag}(1,1,1,1,1,1,-1,-1,-1,-1)$, then, the generators corresponding to the unbroken gauge symmetry commute with this $P_{0}$. The result is
$\sigma_{2}\otimes S_{6},\quad\sigma_{0}\otimes A_{6},\quad\sigma_{2}\otimes I_{6},\quad\sigma_{2}\otimes I_{4},\quad\sigma_{2}\otimes S_{4},$
$\sigma_{0}\otimes A_{4},\quad\sigma_{1}\otimes A_{6},\quad\sigma_{1}\otimes A_{4},\quad\sigma_{3}\otimes A_{6},\quad\sigma_{3}\otimes A_{4}.$ (5)
In this case, the gauge symmetry $SO(20)$ is broken to $SO(12)\times SO(8)$ [6, 18].
If we further choose a $Z_{2}$ parity matrix at $x_{6}=\pi R_{6}$ as
$P_{1}=(I_{11,1},\sigma_{2}\otimes I_{4}),$ (6)
where $I_{4}$ is a $4\times 4$ unit matirx, the unbroken gauge symmetry is $SO(11)\times SU(4)\times U(1)$ and $U(1)$ denotes a subgroup of $SO(8)$, respectively. This is because the generators commutable with $\sigma_{2}\otimes I_{4}$ are $SU(4)$ generators $\sigma_{2}\otimes S_{4},\sigma_{0}\otimes A_{4}$ and a $U(1)$ generator $\sigma_{2}\otimes I_{4}$ in $SO(8)$ generators $\sigma_{2}\otimes S_{4},\sigma_{0}\otimes A_{4},\sigma_{2}\otimes I_{4},\sigma_{1}\otimes A_{4},\sigma_{3}\otimes A_{4}$.
A 6D Dirac $SO(20)$ spinorial representation ${\bf 1024}$ can be decomposed as ${\bf 1024}={\bf 512}_{+}+{\bf 512}_{-}$, where $\pm$ means $Z_{2}$ parity eigenvalues of $\Gamma^{21}$ in $SO(20)$. We introduce only ${\bf 512}_{+}$ to our model. We note that $512_{+}$ is 6D Dirac fermion, which has eight components. $Z_{2}$ parity transformations $x_{i}-x_{6}\to x_{i}+x_{6}$ of 6D Dirac fermion are given
${\bf 512}_{+}(x_{i}-x_{6})=\eta_{i}P_{i}(i\Gamma^{6}){\bf 512}_{+}(x_{i}+x_{6}),$ (7)
where $\eta_{i}$ is a sign ambiguity ±$\pm 1$. Now, we take $(\eta_{0},\eta_{1})=(+,+)$ so that the four components of ${\bf 512}_{+}$ have $+1$ eigenvalues of $i\Gamma^{6}$ and remain to be massless. In the symmetry breaking $SO(20)\to SO(11)\times SU(4)\times U(1)$ by $P_{0}$ and $P_{1}$, ${\bf 512}_{+}$ is decomposed into the following irreducible representations [26, 27].
$\bm{512}_{+}$ $=(\bm{32},\bm{8}_{s})\oplus(\bm{32},\bm{8}_{c})$
$=[(\bm{32},\bm{1}_{2})\oplus(\bm{32},\bm{1}_{-2})\oplus(\bm{32},\bm{6}_{0})]+[(\bm{32},\bm{4}_{-1})\oplus(\bm{32},\bm{\overline{4}}_{1})],$ (8)
where the representations in the parentheses are those of $SO(11)\times SO(8)$ in the first line, $SO(11)\times SU(4)\times U(1)$ in the second line and the number in the subscript implies a $U(1)$ charge. The subscripts $s,c$ of ${\bf 8}$ mean the spinor and conjugated spinor representations of $SO(8)$ [26, 27].
If we assume that a subgroup $Sp(4)$ of $SU(4)$ is a confining gauge group as QCD in the SM [6], the corresponding decomposition of spinors into the irreducible representations of $Sp(4)$ are as follows [26, 27].
$(\bm{32},\bm{8}_{s})=(\bm{32},\bm{1}_{2})\oplus(\bm{32},\bm{1}_{-2})\oplus(\bm{32},\bm{1}_{0})\oplus(\bm{32},\bm{5}_{0}),$ (9)
$(\bm{32},\bm{8}_{c})=(\bm{32},\bm{4}_{-1})\oplus(\bm{32},\bm{4}_{1}).$ (10)
Since the massless spinors should correspond to the $Sp(4)$ singlets, we find three massless $\bm{32}$ spinors, which will provide three generations of quarks and leptons in a model of 5D $SO(11)$ gauge-Higgs grand unification. In the compactification on the orbifold $S^{1}/Z_{2}$ to 4D, we can impose $Z_{2}$ parity such that the left-handed Weyl fermion ${\bf 16}_{L}$ is massless, where three generations of the SM fermions are only realized in the zero mode sector and no other massless fermions are present.
As for the 6D massless gauge field $A_{M}~{}(M=0,1,2,3,5)$ of $SO(20)$ in ${\bf 190}$ dimensional adjoint representation, it is decomposed under $SO(12)\times SO(8)$ and $SO(11)\times SU(4)\times U(1)$ symmetry breaking as [26, 27]
${\bf 190}$ $=({\bf 66},{\bf 1})+({\bf 1},{\bf 28})+({\bf 12},{\bf 8}_{v})~{}(SO(12)\times SO(8)),$ (11)
$=({\bf 55}+{\bf 11},{\bf 1}_{0})+({\bf 1},{\bf 15}_{0}+{\bf 6}_{2}+{\bf 6}_{-2}+{\bf 1}_{0}),$
$\hskip 85.35826pt+({\bf 11}+{\bf 1},{\bf 4}_{-1}+\overline{{\bf 4}}_{1})~{}(SO(11)\times SU(4)\times U(1))$ (12)
${\bf 8}_{v}$ is a vector representation of $SO(8)$.
$(SO(11),SU(4))_{U(1)}$ ($P_{0},P_{1}$)
(${\bf 55}$, ${\bf 1}_{0}$) (+, +)
(${\bf 11}$, ${\bf 1}_{0})$ (+,)(+,-)( + , - )
(${\bf 1}$, ${\bf 15}_{0})$ (+, +)
(${\bf 1}$, ${\bf 6}_{2})$ (+,)(+,-)( + , - )
(${\bf 1}$, ${\bf 6}_{-2})$ (+,)(+,-)( + , - )
(${\bf 1}$, ${\bf 1}_{0})$ (+, +)
(${\bf 11}$, ${\bf 4}_{-1})$ (,+)(-,+)( - , + )
(${\bf 11}$, $\overline{{\bf 4}}_{1})$ (,)(-,-)( - , - )
(${\bf 1}$, ${\bf 4}_{-1})$ (,+)(-,+)( - , + )
(${\bf 1}$, $\overline{{\bf 4}}_{1})$ (,)(-,-)( - , - )
Table 1: $Z_{2}$ parity assignments $(P_{0},P_{1})$ for various components in 6D $SO(20)$ gauge field $A_{M}$.
The $Z_{2}$ parity assignments for the various components in 6D $SO(20)$ gauge field are summarized in Table 1. We can see from this table that the gauge fields of $SO(11)\times SU(4)\times U(1)$ only have $Z_{2}$ parity (+,+)(+,+)( + , + ), which means that they have massless modes. Furthermore, we have two massless mode from the sixth component of the gauge field $A_{6}$, $({\bf 11},\overline{{\bf 4}}),({\bf 1},\overline{{\bf 4}})$ since the $Z_{2}$ parity is totally opposite to those of the 5D gauge field (2). However, these fields are non-singlet in a confining subgroup $Sp(4)$ in $SU(4)$. After all, no massless scalar field comes from $A_{6}$. Note that 5D gauge field $({\bf 1},{\bf 15}_{0})$ also confines and decoupled since no $Sp(4)$ singlet is included in ${\bf 15}$ of $SU(4)$ as can be seen from the decomposition of ${\bf 15}$ of $SU(4)$ into $Sp(4)$ irreducible representations ${\bf 15}={\bf 10}+{\bf 5}$ [26, 27]. To summarize, the massless fields after compactification in the sixth dimension are the 5D $SO(11)\times U(1)$ gauge fields and three spinors in ${\bf 32}$ representation corresponding to three generations of the SM fermions.
In order to obtain 4D theory by the orbifold compactification $S^{1}/Z_{2}$ in the fifth extra space, the following $Z_{2}$ parties are considered.
$A_{\mu}(x_{\mu},x_{i}-x_{5})$ $=P^{\prime}_{i}A_{\mu}(x_{\mu},x_{i}+x_{5})P_{i}^{{}^{\prime}{\dagger}},$ (13)
$A_{5}(x_{\mu},x_{i}-x_{5})$ $=-P^{\prime}_{i}A_{6}(x_{\mu},x_{i}+x_{5})P_{i}^{{}^{\prime}{\dagger}},$ (14)
where $Z_{2}$ matrices $P^{\prime}_{i}~{}(i=0,1)$ at the fixed point $x_{5}=0,\pi R_{5}$ employed in [28]
$P^{\prime}_{0}=I_{10,1},\quad P^{\prime}_{1}=I_{4,7}.$ (15)
In terms of these $Z_{2}$ parities, the gauge symmetry breaking is given by [^1] Here, $SU(4)$ including a confining subgroup $Sp(4)$ is neglected.
$SO(11)\times U(1)$ $\to SO(6)\times SO(4)\times U(1).$ (16)
As for the symmetry breaking to the SM gauge group, see [28].
A 5D $SO(11)$ massless gauge field is decomposed into $SO(6)\times SO(4)$ as follows.  [^2] Here, $U(1)$ charges are neglected.
${\bf 55}=({\bf 15},{\bf 1})+({\bf 6},{\bf 1})+({\bf 6},{\bf 4})+({\bf 1},{\bf 4})+({\bf 1},{\bf 6}).$ (17)
$(SO(6),SO(4))$ $(P^{\prime}_{0},P^{\prime}_{1})$
(${\bf 15}$, ${\bf 1})$ (+, +)
(${\bf 6}$, ${\bf 1})$ (,+)(-,+)( - , + )
(${\bf 6}$, ${\bf 4})$ (+,)(+,-)( + , - )
(${\bf 1}$, ${\bf 4})$ (,)(-,-)( - , - )
(${\bf 1}$, ${\bf 6})$ (+, +)
Table 2: $Z_{2}$ parity assignments $(P^{\prime}_{0},P^{\prime}_{1})$ for the various components in 5D $SO(11)$ gauge field under $SO(6)\times SO(4)$.
The $Z_{2}$ parity assignments for the components in 5D $SO(11)$ gauge field is summarized in Table 2. We can see from the table that the gauge fields of $SO(6)\times SO(4)$ having $Z_{2}$ parity (+,+)(+,+)( + , + ) are massless. On the other hand, $({\bf 1},{\bf 4})$ from $A_{5}$ are found to be massless and identified with the SM Higgs field. As for the fermion ${\bf 32}$, we can obtain three massless ${\bf 16}$ by following [28].
Yukawa coupling can be obtained from the gauge coupling through the fifth component of the gauge field $A_{5}$ in $SO(20)$ theory as follows.
${\bf 512}\cdot{\bf 190}_{H}\cdot{\bf 512}$ $\supset{\bf 32}\cdot{\bf 55}_{H}\cdot{\bf 32}+\cdots({\rm 5D}~{}SO(11))$
$\supset{\bf 16}\cdot{\bf 10}_{H}\cdot{\bf 16}+\cdots,({\rm 4D}~{}SO(10))$ (18)
where the scalar field with subscript $H$ means that it includes the SM Higgs field. Also, the symbol $\supset$ implies that it extracts zero mode terms after compactification. In the first (second) line, the gauge coupling of 6D $SO(20)$ theory is decomposed into 5D $SO(11)$(4D $SO(10)$) symmetry. Since the gauge coupling is universal in flavor space, the fermion mass hierarchy, flavor mixing and CP violation cannot be explained as it stands. Improving these points is very challenging problem and requires nontrivial extensions.
In summary, we have proposed a simple model of family unification based on the grand gauge-Higgs unification. The model is a 6D $SO(20)$ gauge theory with a single fermion in a spinorial representation. Remarkably, three generations of the SM fermions can be derived from a single fermion by assuming a subgroup $Sp(4)$ to be a confining gauge group. The 5D model obtained after compactification of the sixth dimension is a model of $SO(11)\times U(1)$ grand gauge-Higgs unification where the SM Higgs field is unified as the fifth component of the 5D gauge field as well as only three generations of the SM fermions are massless in a fermion sector. Therefore, we believe that our model is the simplest model of family unification which further unifies the gauge fields and the Higgs field.
There are issues to be addressed to discuss phenomenology based on this model. In family unified GHU models, Yukawa coupling constants are universal in flavor space since they are originated from the gauge coupling constant. How the fermion mass hierarchy, flavor mixing and CP violation are realized in our model is a common and important problem to be solved in models of family unification. The issues of gauge coupling unification and proton decay analysis are also important and interesting. In particular, identifying the main decay mode in our model is crucial. Although two step compactifications of 6D and the confining gauge dynamics of $Sp(4)$ were assumed in this paper, a model building based on a compactification on an orbifold $T^{2}/Z_{N}$ and that without any confining gauge group would be more interesting. These are left for our future work.